OpenStudy (anonymous):

Find the sum of the arithmetic sequence. 15, 17, 19, 21, ..., 33

OpenStudy (anonymous):

So just do 15+17+19+21 all the way to 33, it's going up my 2 numbers each time.

OpenStudy (anonymous):

240 ! Thank You.

OpenStudy (anonymous):

any idea how???

OpenStudy (john_es):

You also can use the Gauss' trick, \[\frac{(15+33)10}{2}=240\]

OpenStudy (john_es):

10 is the number of odd numbers you add between 15 and 33, both included.

OpenStudy (anonymous):

initial Input segment value as 15...... Internal common difference as 2.... final Output assigned value as 33 and get the resultant output!!

OpenStudy (anonymous):

xJakester...poor method. And what if I wanted 13 + 15 + 17 + 19 + ......+ 23,117? Are you gonna sit and type all these numbers for a few hours? Correct method is to use the formula for the SUM of an arithmetic series.

OpenStudy (anonymous):

First find out the number of terms in the series. Apply the forum \[Tn = a + (n-1)d\] a = 15, i.e. the first term Tn = nth term, which in this case is 33 d = difference between successive terms = 17-15 = 2 Therefore, 33 = 15 + (n-1) * 2, which gives n = 10 Now apply the formula for the sum of n numbers in an AP \[S = \frac{ n }{ 2 } (2a + (n-1)d)\] This gives you the solution S = 240 (where n = 10)

OpenStudy (anonymous):

I do thank you all for your time, as well as acquiring me with numerous ways to solve this problem :)

OpenStudy (anonymous):

give me a medal..........M waiting

OpenStudy (anonymous):

Jus kiddie........lol